# Tracking Brownian fluid particles in large eddy simulations

### v2

@@ Abstract
In this paper, we establish a numerical method for simulation of wall-bounded incompressible turbulent flows by integrating the technology of random vortex method with the core idea of Large Eddy Simulation (LES). Specifically, we utilize the filtering function in LES, interpreted as spatial averaging, along with the integral representation theorem for parabolic equations, to achieve a closure numerical scheme which may be used for calculating solutions of Navier-Stokes equations. This approach circumvents the challenge associated with handling the non-locally integrable 3-dimensional integral kernel in the random vortex method and facilitates the computation of numerical solutions for flow systems via Monte-Carlo method. Comprehensive numerical simulations are carried out for turbulent and laminar flows in full space and wall-bounded space, considering both two-dimensional and three-dimensional cases, thereby demonstrating the validity and effectiveness of the method.
Key words: large eddy simulation, random vortex method, incompressible fluid flow, turbulence simulation
@@ 1 Introduction
The main goal of the present paper is to propose a numerical scheme for simulating wall-bounded viscous and incompressible fluid flows, based on a stochastic integral representation of solutions to linear parabolic equations (see Section 5 below) to be established in the present paper, together with a key idea from the large eddy simulation (LES, cf. [16], [10] and [5], also cf. the monographs [1] and [9]). The method also applies to ideal models for viscous incompressible fluid flows in the whole space. LES is a numerical method extensively used for simulating turbulent flows (cf. [9] and [20]). In this article, we introduce a natural closure to LES models by using a stochastic formulation of viscous incompressible fluid flows in terms of Brownian fluid particles.
An incompressible viscous flow past a flat plate is modeled with a time dependent velocity
$u(x,t)=(u^{1}(x,t),\cdots,u^{d}(x,t))$
in $D:=\mathbb{R}_{+}^{d}=\{x:x\in\mathbb{R}^{d},x_{d}>0\}$ (where $d=2$ or $3$), so its boundary where $x_{d}=0$ models a solid wall. The fluid density $\rho(x,t)$ and pressure $p(x,t)$ are other fluid flow quantities in this model. Due to the incompressibility of the fluid flow, $\rho(x,t)$ stays as a constant, thus, without losing generality, we may assume that $\rho(x,t)=1$ for all $x$ and $t$. The dynamics of the fluid flow is described by the Navier-Stokes equations
$\frac{\partial}{\partial t}u^{i}+(u\cdot\nabla)u^{i}=\nu\Delta u^{i}-\frac{\partial}{\partial x_{i}}p+F^{i}$ (1)
for $i=1,\ldots,d$, and
$\nabla\cdot u=0$ (2)
@@
If a viscous fluid flow moves along a solid wall with a high speed, then substantial vorticity may be created at the boundary which in turn generates turbulent motion in a thin layer next to the boundary, and may lead to flow separation from the wall. The small-scale energy dissipation in turbulent motion may be described by Kolmogorov’s theory, known as the K41 theory, while the global motion of boundary turbulence may be calculated by means of LES, or other numerical schemes.
The present work aims to propose a method for implementing LES by taking advantage of a stochastic integral representation of the velocity vector field of incompressible viscous fluid flows past a solid wall.
To motivate our approach, let us review the main ideas in the well-known vortex method (cf. [3], [4] and [11]). In the random vortex method, the main fluid dynamical variable is the vorticity, $\omega=\nabla\wedge u$, whose dynamics is described by the vorticity transport equation
$\frac{\partial}{\partial t}\omega^{i}+(u\cdot\nabla)\omega^{i}=\nu\Delta\omega^{i}+(\omega\cdot\nabla)u^{i}+(\nabla\wedge F)^{i}$ (3)
@@
It is appropriate to set up a convention about two-dimensional (2D) fluid flows. By a 2D flow we mean a 3D flow with velocity $u=(u^{1},u^{2},u^{3})$ where $u^{3}=0$, and $u^{1}$, $u^{2}$ depend only on coordinates $x_{1}$, $x_{2}$. Hence, for a 2D flow, $\omega^{1}=0$, $\omega^{2}=0$ identically, $\omega^{3}=\frac{\partial}{\partial x_{1}}u^{2}-\frac{\partial}{\partial x_{2}}u^{1}$, and $\omega$ is identified as a scalar function – its third component $\omega^{3}$, unless otherwise indicated. In particular the vorticity stretching term $(\omega\cdot\nabla)u$ vanishes identically for a 2D flow, and the vorticity equation becomes
$\frac{\partial}{\partial t}\omega+(u\cdot\nabla)\omega=\nu\Delta\omega+\nabla\wedge F$
@@
Two equations (5, 6) consist of a closed McKean-Valsov type stochastic differential equations, which in turn may be used in the design of numerical schemes for calculating the flow velocity. It is possible to formulate the random vortex scheme for three-dimensional viscous flows, which, however, involves an integral kernel which is not locally integrable, leading to difficulties in numerical computation of 3D flows. See more details and related work in [7] and [14].
It is also possible to derive a random vortex formulation for flows constrained on a finite region and for flows past a solid wall. Although the random vortex method is mathematically elegant, it is effective only for certain types of flows and is often computationally expensive. For simulations of turbulent flows, schemes for calculating solutions of turbulent flows from their dynamical equations at separate space scales shall be advocated. At small scale, the universality principle of turbulence implies that the statistical properties of turbulence, such as energy dissipation and the structure of velocity gradients, become independent of the large-scale features of the flow and are governed by universal scaling laws. This motivates us to develop a new random vortex method that focuses more on relatively large scale flows.
The main contribution of the present work is that the integral representation of the fluid flow in terms of Brownian fluid particles gives rise to a natural closure of the averaged Navier-Stokes equations without further modelling the stress tensor $\tau$.
@@
The remainder of the paper is organized as follows. In Section 2, the theoretical foundation of the random LES, proposed in this work, will be established including the details of the formulation of incompressible viscous fluid flows in terms of Brownian fluid particles. In Section 3, the random LES flows past a wall is established. The random LES are formulated in terms of stochastic differential equations involving the distribution of Brownian fluid particles. Various numerical experiments based on random LES are reported in Section 4. Finally, in an appendix, Section 5, we include the necessary results on the functional integral representation for solutions of linear parabolic systems, used in the formulation of the random LES.
@@ 2 Theoretical aspect of random LES
Let us first reformulate the fluid dynamics in terms of Brownian fluid particles, a device firstly put forward by G. Taylor [19]. Although it has not been established rigorously yet, we will nevertheless assume that the velocity $u(x,t)$ is bounded and smooth in $D\times[0,T)$ (for some $T>0$) and at least $C^{1}$ up to the boundary $\partial D$. As a consequence, $p(x,t)$ shares the same regularity as that of $u(x,t)$. The dynamics of the flow is determined by the Navier-Stokes equations (1, 2) together with the non-slip boundary condition. We introduce the theoretical framework of random LES.
@@ 2.1 Computation of the pressure gradient
The method we are going to propose, which combines the idea of Brownian fluid particles with an idea from LES, is based on a functional integral representation of solutions to linear parabolic equations in terms of distributions of Brownian fluid particles, cf. Section 5. Unlike the random vortex approach, although still utilizing the distribution of the Brownian fluid particles, we calculate the velocity directly instead of dealing with the vorticity equation. Hence we need to resolve the pressure gradient $\nabla p$ first.
@@
$\frac{\partial}{\partial\boldsymbol{n}}p=\nu\left(\frac{\partial}{\partial x_{1}}\frac{\partial u^{1}}{\partial x_{3}}+\frac{\partial}{\partial x_{2}}\frac{\partial u^{2}}{\partial x_{3}}\right)-F^{3}\quad\textrm{ on }\partial D$ (8)
@@
for $d\geq 3$, where $s_{d-1}$ is the area of $d-1$ dimensional sphere, $s_{1}=2\pi$ and $s_{2}=4\pi$.
@@
$\;+\int_{\mathbb{R}^{d-1}}H_{d}^{+}(x,(y_{1},\cdots,y_{d-1},0))\left.\frac{\partial p}{\partial y_{d}}\right|_{y_{d}=0}(y_{1},\cdots,y_{d-1},0)\textrm{d}y_{1}\cdots\textrm{d}y_{d-1}$ (13)
@@
$K_{3}^{+}((x_{1},x_{2},x_{3}),(y_{1},y_{2},0))=\frac{1}{2\pi}\frac{\left(x_{1}-y_{1},x_{2}-y_{2},x_{3}\right)}{\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{3}{2}}}$ (18)
@@
$\;+\nu\int_{\mathbb{R}}\frac{\partial K_{2}^{+}}{\partial y_{1}}(x,(y_{1},0))\frac{\partial u^{1}}{\partial y_{2}}(y_{1},0)\textrm{d}y_{1}$ (19)
@@
$\;+\nu\int_{\mathbb{R}^{1}}\frac{\partial K_{2}^{+}}{\partial y_{1}}(x,(y_{1},0))\frac{\partial u^{1}}{\partial y_{2}}(y_{1},0)\textrm{d}y_{1}$
@@
$\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,y)=\frac{\left(2|x_{1}-y_{1}|^{2}-|x_{2}-y_{2}|^{2}-|x_{3}|^{2},3(x_{2}-y_{2})(x_{1}-y_{1}),3x_{3}(x_{1}-y_{1})\right)}{2\pi\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{5}{2}}}$ (22)
@@
$\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,y)=\frac{\left(3(x_{1}-y_{1})(x_{2}-y_{2}),2|x_{2}-y_{2}|^{2}-|x_{1}-y_{1}|^{2}-|x_{3}|^{2},3x_{3}(x_{2}-y_{2})\right)}{2\pi\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{5}{2}}}$ (23)
@@
and $s_{d-1}$ is the area of the $d-1$ dimensional unit sphere, $s_{1}=2\pi$ and $s_{2}=4\pi$.

### v4

# Tracking Brownian fluid particles in large-eddy simulations

@@ Abstract
In this paper, we propose an approach for simulating wall-bounded incompressible turbulent flows by integrating the technology of random vortex method with the core principles of large-eddy simulations (LES). In particular, we employ the filtering function, interpreted as a spatial averaging operator, together with the integral representation theorem for parabolic equations, to construct a closed numerical scheme suitable for computing solutions to the Navier-Stokes equations. This framework numerically overcomes the difficulties associated with the non-locally integrable three-dimensional kernel inherent in the random vortex method, enabling efficient computation of flow fields via the Monte Carlo method. Several numerical experiments are presented for both laminar and turbulent flows in wall-bounded domains, to thereby reveal the underlying flow mechanisms near the wall boundary. The experimental results and systematic comparisons with alternative numerical approaches consistently demonstrate that the proposed method is numerically stable, possesses low theoretical complexity, and achieves acceptable computational efficiency.
Key words: large-eddy simulation, random vortex method, incompressible fluid flow, turbulence simulation
@@ 1 Introduction
In the present paper, we propose a novel scheme for implementing large-eddy simulations (LES, cf. [25, 17, 8], also cf. the monographs [1, 16]) for wall-bounded viscous and incompressible fluid flows, based on a stochastic integral representation of solutions to linear parabolic equations (see Appendix A below) to be established in the present paper. LES is a numerical method extensively used for simulating turbulent flows (cf. [16, 29, 15]). In this article, we introduce a natural closure to LES models by using a stochastic formulation of viscous incompressible fluid flows in terms of Brownian fluid particles.
An incompressible viscous flow past a flat plate is modeled with a time-dependent velocity
$u(x,t)=(u^{1}(x,t),\cdots,u^{d}(x,t)),$
in $D:=\mathbb{R}_{+}^{d}=\{x:x\in\mathbb{R}^{d},x_{d}>0\}$ (where $d=2$ or $3$), so its boundary where $x_{d}=0$ models a solid wall. The fluid density $\rho(x,t)$ and pressure $p(x,t)$ are other fluid flow quantities in this model. Due to incompressibility of the fluid flow, $\rho(x,t)$ is constant, thus, without losing generality, we may assume that $\rho(x,t)=1$ for all $x$ and $t$. The dynamics of the fluid flow is described by the Navier-Stokes equations
$\frac{\partial}{\partial t}u^{i}+(u\cdot\nabla)u^{i}=\nu\Delta u^{i}-\frac{\partial}{\partial x_{i}}p+F^{i},$ (1)
, for $i=1,\ldots,d$, and
$\nabla\cdot u=0,$ (2)
@@
The Navier-Stokes equations may be formulated in various weak solution forms, which allows implementing the finite difference method of solving numerically their solutions. To this end, the mild solution approach seems useful. The simplest mild solution formulation is achieved by using the heat kernel. Let $h(x,t,y)$ be the heat kernel of the heat operator $\nu\Delta-{\partial\over\partial t}$. For any given $t>0$, considering $f(s)=\int u(y,t-s)h(x,s,y)\textrm{d}y$. Then $f(0)=u(x,t)$, $f(t)=u_{0}(x)$, and
$f(t)-f(0)$ $=\int_{0}^{t}f^{\prime}(s)\textrm{d}s$
$=\int_{0}^{t}\int_{D}u(y,t-s){\partial\over\partial s}h(x,s,y)\textrm{d}y\textrm{d}s-\int_{0}^{t}\int_{D}{\partial\over\partial s}u(y,t-s)h(x,s,y)\textrm{d}y\textrm{d}s$
$=\int_{0}^{t}\int_{D}\left(\nu\Delta-{\partial\over\partial s}\right)u(y,t-s)h(x,s,y)\textrm{d}y\textrm{d}s$
$=\int_{0}^{t}\int_{D}\left(u\cdot\nabla u+\nabla p-F\right)(y,t-s)h(x,s,y)\textrm{d}y\textrm{d}s,$ (3)
which gives rise to the following implicit representation of the velocity
$u(x,t)=u_{0}(x)-\int_{0}^{t}\int_{D}\left(u\cdot\nabla u+\nabla p-F\right)(y,s)h(x,t-s,y)\textrm{d}y\textrm{d}s,$ (4)
for every $t>0$ and $x\in D$. The gradient $\nabla u$ appearing in the integral on the right-hand side can be removed by using integration by parts due to the incompessibility of $u$, and
$u^{i}(x,t)-u^{i}_{0}(x)=$
$\int_{0}^{t}\int_{D}\left(u^{i}(y,s)u(y,s)\cdot\nabla_{y}\ln{h}(x,t-s,y)-{\partial\over\partial y_{i}}p(y,s)-F^{i}(y,s)\right)h(x,t-s,y)\textrm{d}y\textrm{d}s,$ (5)
for $i=1,2,3$. This mild solution setting may be used to implement finite difference method and calculate numerical solutions. However, this method requires still, though not explicitly involving the gradient of $u$ in the formulation, the iteration of $\nabla u$ in order to calculate the pressure gradient $\nabla p$. This means that we need to use finite difference method to calculate $\nabla u$, which brings the risk of numerical explosion. There is another technical obstacle in this mild solution formulation, that is, the time of both singular integral kernel $\nabla\ln{h}$ and h(x,ts,y)h(x,t-s,y) are reversed, which implies that we need to save and use the velocity of every time during the iterative computations, causing a very long computing time. The mild solution formulation is widely used in the study of stochastic partial differential equations, it seems not suitable for directly numerical simulation of turbulent flows, further detailed explanation and comparison on this aspect shall be addressed in subsection 4.3.2.
To overcome this difficulty of the previous mild solution formulation one may try to follow the main idea in this formulation by absorbing the quadratic term $u\cdot\nabla u$ into the Laplacian term and make use of the fundamental solution associated with the forward parabolic operator $\nu\Delta+u\cdot\nabla-{\partial\over\partial t}$ in place of the Gaussian heat kernel $h(x,t,y)$, which is the transition probability of the diffusion with velocity $u(x,t)$. This approach can be traced back to the original work by Chorin [5] which gives the life of the random vortex method, which leads to the increasing attention paid to vortex [3, 32, 33, 2].
Let us review the main ideas in this approach for more details, cf. [5, 7] and [18]. In the random vortex method, the main fluid dynamic variable is taken to be the vorticity, $\omega=\nabla\wedge u$, whose dynamics is described by the vorticity transport equation
$\frac{\partial}{\partial t}\omega^{i}+(u\cdot\nabla)\omega^{i}=\nu\Delta\omega^{i}+(\omega\cdot\nabla)u^{i}+(\nabla\wedge F)^{i},$ (6)
@@
It is appropriate to set up a convention about two-dimensional (2D) fluid flows. By a 2D flow we mean a three-dimensional (3D) flow with velocity $u=(u^{1},u^{2},u^{3})$ where $u^{3}=0$, and $u^{1}$, $u^{2}$ depend only on coordinates $x_{1}$, $x_{2}$. Hence, for a 2D flow, $\omega^{1}=0$, $\omega^{2}=0$ identically, $\omega^{3}=\frac{\partial}{\partial x_{1}}u^{2}-\frac{\partial}{\partial x_{2}}u^{1}$, and $\omega$ is identified with a scalar function – its third component $\omega^{3}$, unless otherwise indicated. In particular the vorticity stretching term $(\omega\cdot\nabla)u$ vanishes identically for a 2D flow, and the vorticity equation becomes
$\frac{\partial}{\partial t}\omega+(u\cdot\nabla)\omega=\nu\Delta\omega+\nabla\wedge F,$ (7)
@@
Two equations (9, 10) consist of a closed McKean-Valsov type stochastic differential equations, which in turn may be used in the design of numerical schemes for calculating the flow velocity. It is also feasible to derive a random vortex formulation for flows constrained on a finite region and for flows past a solid wall. Although the random vortex method is mathematically elegant, it is effective only for certain types of flows and is sometimes computationally expensive.
It is possible to formulate the random vortex scheme for three-dimensional viscous flows, which, however, involves an integral kernel which is not locally integrable, leading to difficulties in numerical computation of 3-dimensional flows. This poses a significant obstacle to the application of the random vortex method in the simulation of three-dimensional flows. In comparison, the random vortex method is more suitable for handling two-dimensional cases because the equations are simpler, and the integral kernel can be processed using a smoothing technique. See more details and related work in [11] and [23]. The significance, at least formally, the quadratic term $u\cdot\nabla u$ is hidden in the previous integral representation of the velocity in terms of the Brownian particle diffusion, which appears linear in $u$.
If a viscous fluid flow moves along a solid wall with a high speed, then substantial vorticity may be created at the boundary which in turn generates turbulent motion in a thin layer next to the boundary, and may lead to flow separation from the wall. The small-scale energy dissipation in turbulent motion may be described by Kolmogorov’s theory, known as the K41 theory, while the global motion of boundary turbulence may be calculated by means of LES, or other numerical schemes. For simulations of turbulent flows, methods such as finite difference and finite element require extremely fine mesh resolution, which places high demands on computational power, therefore, schemes for calculating solutions of turbulent flows from their dynamical equations at separate space scales shall be advocated. LES method (cf. [21, 25]), however, uses a filtering function to separate large eddies from small eddies: the large eddies are computed numerically, while the small eddies are modeled using standardized models. On a small scale, the universality principle of turbulence implies that the statistical properties of turbulence, such as energy dissipation and the structure of velocity gradients, become independent of large-scale features of the flow and are governed by universal scaling laws [25, 10, 13, 14]. The integration of LES and machine learning discovers superior closure models and algorithms in a data-driven manner, significantly enhancing computational efficiency while ensuring accuracy [31, 30]. LES approach greatly reduces computational cost and has become the primary method for turbulence simulation today. Compared to the random vortex method, LES suffers from significant grid dependency, inherent numerical dissipation that smears small-scale vortices, and greater complexity in handling moving boundaries. This motivates us to develop a new random vortex method that focuses more on relatively large-scale flows.
To address the challenges faced by the random vortex method in three-dimensional simulations and to incorporate the concept of large-eddy modeling, we aim to propose a random calculation method for implementing LES. The contribution of this paper lies in combining LES with the stochastic integral representation theorem to propose an explicit, forward-in-time computational method for the Navier-Stokes equations. This random LES approach can be adopted to simulate three-dimensional wall-bounded turbulence using the Monte Carlo method. According to our experimental results, our method is tolerant of mesh resolution, does not require high computational power, remains numerically stable without blow-up, and has acceptable computation time. Moreover, the method supports repeatability, allowing the simulation of fluid behavior under potential distributions. Our experimental results reveal some of the flow mechanisms at the wall boundary. Compared with the stochastic vortex method, our approach overcomes the challenge of handling three-dimensional fluid systems. Compared with traditional LES, it closes the Navier-Stokes equations without further modelling the stress tensor and imposes less stringent demands near the wall.
@@
The remainder of the paper is organized as follows. In Section 2, the theoretical foundation of the random LES, proposed in this work, will be established including the details of the formulation of incompressible viscous fluid flows in terms of Brownian fluid particles. In Section 3, the random LES flows past a wall is established. The random LES are formulated in terms of stochastic differential equations involving the distribution of Brownian fluid particles. Various numerical experiments based on random LES, comparisons, and analysis are reported in Section 4. Finally, in Appendix, we include the necessary results on the functional integral representation for solutions of linear parabolic systems, used in the formulation of the random LES.
@@ 2 Theoretical aspect of random LES
Let us first reformulate the fluid dynamics in terms of Brownian fluid particles, following G.Taylor [28]. Although it has not been established rigorously yet, we will nevertheless assume that the velocity $u(x,t)$ is bounded and smooth in $D\times[0,T)$ (for some $T>0$) and at least $C^{1}$ up to the boundary $\partial D$. As a consequence, $p(x,t)$ shares the same regularity as that of $u(x,t)$. The dynamics of the flow is determined by the Navier-Stokes equations (1, 2) together with the non-slip boundary condition. We introduce the theoretical framework of random LES.
@@ 2.1 Computation of the pressure gradient
The method we are going to propose, which combines the idea of Brownian fluid particles with an idea from LES, is based on a functional integral representation of solutions to linear parabolic equations in terms of distributions of Brownian fluid particles, cf. Appendix A. Unlike the random vortex approach, although still utilizing the distribution of the Brownian fluid particles, we calculate the velocity directly instead of dealing with the vorticity equation. Hence we need to resolve the pressure gradient $\nabla p$ first.
@@
$\frac{\partial}{\partial\boldsymbol{n}}p=\nu\left(\frac{\partial}{\partial x_{1}}\frac{\partial u^{1}}{\partial x_{3}}+\frac{\partial}{\partial x_{2}}\frac{\partial u^{2}}{\partial x_{3}}\right)-F^{3}\quad\textrm{ on }\partial D,$ (12)
@@
for $d\geq 3$, where $s_{d-1}$ is the area of d1d-1 dimensional sphere, $s_{1}=2\pi$ and $s_{2}=4\pi$.
@@
$\;+\int_{\mathbb{R}^{d-1}}H_{d}^{+}(x,(y_{1},\cdots,y_{d-1},0))\left.\frac{\partial p}{\partial y_{d}}\right|_{y_{d}=0}(y_{1},\cdots,y_{d-1},0)\textrm{d}y_{1}\cdots\textrm{d}y_{d-1},$ (17)
@@
$K_{3}^{+}((x_{1},x_{2},x_{3}),(y_{1},y_{2},0))=\frac{1}{2\pi}\frac{\left(x_{1}-y_{1},x_{2}-y_{2},x_{3}\right)}{\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{3}{2}}},$ (22)
@@
$\;+\nu\int_{\mathbb{R}^{1}}\frac{\partial K_{2}^{+}}{\partial y_{1}}(x,(y_{1},0))\frac{\partial u^{1}}{\partial y_{2}}(y_{1},0)\textrm{d}y_{1},$
@@
$\frac{\partial K_{3}^{+}}{\partial y_{1}}(x,y)=\frac{\left(2|x_{1}-y_{1}|^{2}-|x_{2}-y_{2}|^{2}-|x_{3}|^{2},3(x_{2}-y_{2})(x_{1}-y_{1}),3x_{3}(x_{1}-y_{1})\right)}{2\pi\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{5}{2}}},$ (26)
@@
$\frac{\partial K_{3}^{+}}{\partial y_{2}}(x,y)=\frac{\left(3(x_{1}-y_{1})(x_{2}-y_{2}),2|x_{2}-y_{2}|^{2}-|x_{1}-y_{1}|^{2}-|x_{3}|^{2},3x_{3}(x_{2}-y_{2})\right)}{2\pi\left(|x_{1}-y_{1}|^{2}+|x_{2}-y_{2}|^{2}+|x_{3}|^{2}\right)^{\frac{5}{2}}}.$ (27)
@@
and $s_{d-1}$ is the area of the d1d-1 dimensional unit sphere, $s_{1}=2\pi$ and $s_{2}=4\pi$.